A chemist wants to make 44 mL of a 16% acid solution by mixing a 10% acid solution and an 18% acid solution. How many milliliters of each solution should the chemist use?
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

---------percent ---------------- quantity
Acid solution I 10 ---------------- x ml
Acid solution II 18 ------ 44 - x ml
Mixture 16.00% ---------------- 44

10x +18 (44-x ) = 16 * 44

10x+792 - 18 x = 704
10x -18x = 704 - 792
-8 x = -88
/ -8
x = 11 ml 10.00% Acid solution I
33 ml 18.00% Acid solution II


Answer by stanbon(75887)

A chemist wants to make 44 mL of a 16% acid solution by mixing a 10% acid solution and an 18% acid solution. How many milliliters of each solution should the chemist use?
-----
Equation:
acid + acid = acid
0.10x + 0.18(44-x) = 0.16*44
-----
10x + 18*44 - 18x = 16*44
-8x = -2*44
x = 11 mL (amt. of 10% solution needed)
44-x = 33 mL (amt. of 18% solution needed)
===================
Cheers,
Stan H.
=====================